Last year, X Company sold 66,100 units of its only product for $18.00 each. Total costs were as follows:

Cost of goods sold
   Variable $442,209
   Fixed 120,302
Selling and administrative
   Variable $78,659
   Fixed 66,100


At the end of the year, a company offered to buy 4,100 units of the product but only for $12.00 each. X Company had the capacity to produce the additional units, and even though there would have been no additional selling and administrative costs, it rejected the offer.

5. If X Company had accepted the special order, firm profits would have increased by

6. X Company's production manager thought that the special order units might have required additional direct material and direct labor costs per unit of $0.77 and $0.29, respectively, and the rental of special equipment for $1,000. The combined effect of these changes would have been to reduce the special order profit by

7. X Company's marketing manager felt that in order for regular sales to continue at 66,100 units this year, the company would have had to reduce the regular selling price this year to $17.52. The result of this price reduction would have been to reduce X Company's profits this year by
Answer #1

5)

S.P per unit 18.00
Variable COGS (442209/66100) 6.69
Variable Selling & Distribution (78659/66100) 1.19
Contribution Per unit 10.12
Since the X company had the capacity to produce 4100 units but did not accept the order,
it will result in loss of = 12 - 6.69 - 1.19 = 4.12 per unit
In total the contribution lost for not accepting the order will be(4.12*4100) 16892
Hence if X company had accepted the special order firms profits would have increased by $16892

6)

Increase in cost per unit = 0.77 + 0.29 = 1.06

Total increase in cost = (4100*1.06) + 1000 = 5346

Hence the combined effect of the changes will reduce the special order profit by $5346

7)

Earlier contribution per unit = 10.12

Revised contribution per unit = 17.52 - 6.69 - 1.19 = 9.64

Hence, the result of this price reduction would have been to reduce X Company's profits this year by

(10.12 - 9.64)*66100 units = $31728